OpenStudy (anonymous):

How to convert 8.16666666 into a fraction

OpenStudy (anonymous):

Please help!

OpenStudy (doc.brown):

\[\frac{8.16666666}{1}=\frac{81.6666666}{10}=\frac{816.666666}{100}=\frac{8166.66666}{1000}etc\]\[\frac{816666666}{100000000}\] And you're done

OpenStudy (zarkon):

I assume you mean \(8.1\overline{6}\) let \(x=8.1\overline{6}\). then \(10x=81.\overline{6}\) \[10x-x=81.\overline{6}-8.1\overline{6}\] \[9x=73.5\] \[x=\frac{73.5}{9}=\frac{735}{90}=\frac{49}{6}\]

OpenStudy (doc.brown):

That's the coolest thing I have seen today. Thanks @Zarkon !